For P 100, I can buy 4 chicken pies and 12 banana cakes, or 5 banana cakes and 10 chicken pies. How much does one of each kind cost?
Answer by RAY100(1637)

(1) 100=4C +12B,,,(divide by 4),,,25=1C+3B,,(25-3B) = C
.
(2) 100=10C+5B,,,(divide by 5),,,20=2C+1B
.
subst (1) in (2)
.
20 =2(25-3B) +B
20 = 50 -6B+B
-30= -5B
6=B
.
from (1) C=25-3B
C=25-3(6)
C=7
.
check
(1) 100=4(7)+12(6)= 28+72 = 100,,,,ok
(2) 100 =10(7) +5(6) = 70+30=100,,,ok